# Probing the axial symmetry of gamma-ray burst jets using afterglow polarimetry

### v1

@@ 5 Case Studies
Optical linear polarisation measurements have been conducted for many GRB afterglows, typically several hours to a few days after the prompt gamma-ray emission (Covino & Gotz 2016, and references therein). This timeframe corresponds to when a jet break is expected to occur. The observed linear polarisation degree or its upper limits are generally low ($\;\buildrel<\over{\sim}\;$ several percent). Such low polarisation levels suggest that in shock-generated random magnetic fields, neither the component parallel nor perpendicular to the shock normal dominates entirely (e.g., Gill & Granot 2020), though the fields are not fully isotropic. Our model, which assumes a low local polarisation degree parameter $P_{0}$, should be applicable to modelling afterglow polarisation measurements.
@@ 5.1 GRB 020813
A bright long-duration burst ($>125$ s) detected by HETE-2, located at a redshift of $z=1.255$ (Barth et al., 2003), was studied. Polarimetry data obtained with VLT and Keck (Gorosabel et al., 2004) was collected around a jet break, occurring at $\sim 13$ hours post-burst (Covino et al. 2003). The data is most consistently described by an almost constant degree of linear polarisation at the $\sim 1\%$ level, along with a stable position angle. While a slow evolution of the position angle $\Delta\phi\sim$ a few tens of degrees cannot be ruled out, the observations do not support a sudden 90 degree rotation of the position angle at the jet break, which is a signature of top-hat axisymmetric jets. Given the sparsity of polarisation data points, a single-peaked polarisation curve remains consistent with the observations. The evolution of both the polarisation degree and position angle can explained by structured jets, whether axisymmetric or not. If a gradual rotation of the position angle were confirmed, it would indicate a structured, non-axisymmetric jet.
@@ 5.2 GRB 091018
Linear polarisation monitoring with VLT (Wiersema et al., 2012) revealed a polarisation curve inconsistent with a single-peaked profile, and instead favouring a two-peak structure with a trough around the suspected jet break. In top-hat jet models, both axisymmetric and non-axisymmetric, the second peak is expected to have a higher amplitude than the first. Although some data points support this trend, significant scatter suggests the presence of a deep, short dip within the second peak. The PPA remains stable and nearly constant during the first peak, but shifts to a higher overall value in the second peak, with significant scatter around the 90-degree rotation. These variations in both the polarisation degree and position angle during the second peak are not fully explained even by non-axisymmetric jets (i.e., elliptical jet heads). As noted by Wiersema et al. 2012, if the scatter in the second peak is caused by an additional component (e.g., a bright patch within the jet or microlensing effects), the presence of a two-peak profile, combined with a nearly constant PPA in the first peak, suggests either a top-hat jet or a view of the core of a structured jet. For a non-axisymmetric jet, the line of sight must be positioned near either the semi-major or semi-minor axis of the jet head. Additionally, the shallow post-break decay indices may imply that the jet core is encased by an outer region with a shallow angular profile.
@@ 5.3 GRB 121024A
A long burst with $T_{90}=69$ sec was detected by Swift, with a redshift of $z=2.298$ determined shortly afterwards. The X-ray light curve follows a three-segment power-law evolution, with the final break at $t\sim 10$ hours coinciding with a break in the optical light curve (Wiersema et al., 2014). The monochromatic nature of this break suggests a jet break origin. However, the post-break decay indices remain relatively shallow, at 1.67 in the X-rays and 1.25 in the optical. VLT polarimetry measurements taken 3–6 hours after the burst show linear polarisation degrees ranging from 3–5%, with a stable position angle. However, observations from the following night (after the jet break) reveal a significant shift in the position angle, consistent with a 90 degree rotation (Wiersema et al., 2014). These polarisation properties also support the same jet structure interpretations as for GRB 091018.
@@ 5.4 GRB 210610B
A long-duration burst at $z=1.134$ was detected by Fermi ($T_{90}=69$ sec) and Swift ($T_{90}=55$ sec). The optical and X-ray light curves are modelled using a broken power-law. The optical light curve remains initially flat until the break at $\sim 7.8$ hours, after which it steepens to a decay index of 1.85. The X-ray light curve appears to undergo a simultaneous break and shares the same post break decay index, thought its pre-break decay is faster than that of the optical counterpart (Agüí Fernández et al., 2024). Polarimetric observations were conducted with the Calar Alto Telescope and the VLT. The first measurements, taken at around 2.9 hours, revealed a relatively high polarisation degree of $\sim 4.5\%$. During the second observation period (5.8–6.7 hours), which is close to the break time, the polarisation level dropped significantly to 0.18–0.6%. However, after the break, the polarisation degree increased again to $\sim 2\%$ (30.4–30.6 hours). Interestingly, the PPA rotates by 54 degrees between the pre- and post-break measurements, suggesting a significant shift that deviates notably from the 90 degree rotation expected for axisymmetric jets.
This event is a strong candidate for modelling with a non-axisymmetric jet. Here, we assume top-hat, non-axisymmetric jets for our analysis. While a structured, non-axisymmetric jet could also explain the event if the line of sight intersects the jet core, the constraint on the jet head’s eccentricity remain similar. Given the large deviation of $|\Delta\phi-\pi/2|=36$ degrees, and the distinct trough structure $R_{pd}\geq 10$, we performed 5000 random samplings. For a lower eccentricity ($e=0.3$), most cases (87%) yield $R_{pd}>10$. However, virtually none satisfy $|\Delta\phi-\pi/2|>$ 36 degrees. For $e=0.75$, the fraction of cases meeting the first, second and both conditions are 64%, 35% and 21%, respectively. The lines of sight that satisfy both conditions are characterised by intermediate azimuthal angles: $1.50.4<\chi<$ 1.5 and a relatively central region: $\xi<0.7$. These findings suggest this event could be explained within a non-axisymmetric jet framework. However, there are caveats in the jet-break modelling. 1) The flat optical light curve before the break suggests significant energy injection into the forward shock at early times. Since ejecta from the central engine can carry large-scale magnetic fields (Mundell et al., 2013) unlike forward shocks, which generates highly tangled magnetic fields, the polarisation signals before the break may be significantly influenced by emission from refreshed shocks (energy injection). 2) the post-break decay index of the optical afterglow ($1.85\pm 0.04$) is close to, but still shallower than, the expected value ($\sim 2$) for the standard fireball model (Agüí Fernández et al., 2024).
@@ 6 Conclusions
We have investigated the afterglow polarisation signals around a jet break for non-axisymmetric jets. For jets with top-hat luminosity density distributions, the polarisation degree curves exhibits two peaks in general, even when the jet head is elliptical. However, a complete polarisation drop between the peaks is generally absent. Unlike axisymmetric jets, where the polarisation position angle (PPA) undergoes a sudden jump at the trough, the PPA in non-axisymmetric jets evolves smoothly. Additionally, the change in PPA between the two peaks can deviate significantly from 90 degrees. In our random line-of-sight sampling, 50% of the cases show deviations of more than 18 degrees, and 10% exceed 31 degrees for $e=0.6$.
@@
As a simple toy model for a structured jet, we considered a case where the luminosity density in the jet’s outer region follows a power-law profile. While this simplified model should captures the key features of polarisation signals in structured jets, the angular dependence of luminosity density in a more realistic scenario arises from the lateral energy distribution - where the jet core has a higher energy per solid angle, and the outer regions gradually contain less energy per solid angle. This also induces the angular variations in the Lorentz factor, with the core having a higher Lorentz factor and progressively decreasing toward the outer regions. As a result, the visible region is no longer circular symmetric, toward the jet centre, the relativistic beaming is stronger and luminosity contribution is larger (e.g., Figure 2 in Beniamini et al. 2022). The impact of this deformed visible region will be explored in a future paper.
When constructing the polarisation degree light curves shown in Figs 2, 6 and 7, we assumed the forward shock dynamics $\Gamma\propto t^{-3/8}$ for the ISM. In the wind medium, the curve shapes are slightly modified due to a different forward shock dynamics $\Gamma\propto t^{-1/4}$. However, the overall morphology (e.g. the presence of two peaks or a single peak) remains unchanged and the discussion on $\Delta\phi$ should be similar.
If the jet head has a highly eccentric shape, it can be characterised by two angular scales: $\theta_{j}$ (the opening angle along the semi-major axis) and $\theta_{j2}$ (along the semi-minor axis). As long as $\Gamma>1/\theta_{j2}$, the flux follows the typical blast wave emission evolution, decaying approximately as $t^{-1}$. When $\Gamma\propto t^{-3/8}$ drops below this first critical threshold, photons begin to be emitted over a larger solid angle, increasing from $\theta_{j}\theta_{j2}$ to $\Gamma^{-1}\theta_{j}$, causing the flux to decay as $t^{-11/8}$. As $\Gamma$ continues to decreases and falls below the 2nd threshold $1/\theta_{j}$, the emission solid angle expands further to $\Gamma^{-2}$, leading to a steeper decay of $t^{-7/4}$. Since these breaks arise from geometrical effects, they are expected to be monochromatic, similar to the usual jet break. The two jet break times $t_{j}$ and $t_{j2}$, with $t_{j2}<t_{j}$, satisfy the relation
@@
Although our case study section focused on optical afterglow observations, the jet break is a geometrical effect, meaning our results are applicable across different wavelengths (e.g., X-ray, radio). The number of events with measured radio polarisation has recently increased (Corsi et al. 2018; Laskar et al. 2019; Urata et al. 2023). Significant advancement in X-ray polarimetry in recent years (e.g., Astrosat, IXPE, POLAR, XPoSat) suggest that future X-ray satellites may be capable of detecting and analysing polarisation signals days after GRBs. Since both X-ray and optical bands are expected to lie above the typical frequency of the forward shock emission around jet breaks, the relativistic limb-brightening effect should be pronounced in both bands (Granot et al., 1999).
We adopted a conventional scenario in which magnetic fields are amplified by plasma kinetic instabilities at the shock, resulting in a coherence length scale comparable to the plasma skin depth scale (Medvedev & Loeb, 1999). Recently, large-scale turbulent magnetic fields driven by magnetohydrodynamic instabilities have also been proposed (e.g., Kuwata et al. 2023). In this scenario, the coherence length scale is comparable to the thickness of the blast wave. In the isotropic field case, both the polarisation degree and the PPA change randomly and continuously over time, resembling the behaviour described in the classical patch shell model (e.g., Gruzinov & Waxman 1999). By incorporating magnetic field anisotropy and the observer’s viewing angle, this model can also account for a variety of polarisation behaviour in afterglows. Notably, in this framework, the degree of radio polarisation can exceed that in the optical band (Kuwata et al., 2024). Simultaneous polarimetric observations across multiple wavelengths will be crucial for constraining the magnetic field amplification mechanism and improving our understanding of polarisation signals in afterglows.
@@ Acknowledgements
This research was supported by the Sciences and Technology Facilities Research Council (STFC).

### v2

@@ 5 Case Studies
[…]
@@ 5.1 GRB 020813
Figure 8: Sample fit for GRB020813 (red solid line). Structured jet model parameters: $e=0.6$, $\theta_{c}/\theta_{j}=0.13$, $\xi=3(\theta_{c}/\theta_{j})$, $\chi=0.8$, and $P_{0}=0.15$. Optical polarimetry data from (Gorosabel et al., 2004) are shown, where squares (pre-break) and circles represent Keck and VLT measurements, respectively. The time axis is normalised to the break $t_{j}=0.56\pm 0.21$ days detected in the V-band light curve, with the break-time uncertainty indicated by the shaded region.
A bright long-duration burst (>125>125 s) detected by HETE-2, located at a redshift of $z=1.255$ (Barth et al., 2003), was studied. Polarimetry data obtained with Keck and VLT and (Gorosabel et al., 2004) was collected around a jet break, occurring at $\sim 13$ hours post-burst (Covino et al. 2003). As shown in Figure 8, the data is most consistently described by an almost constant degree of linear polarisation at the $\sim 1\%$ level, along with a stable position angle. While a slow evolution of the position angle $\Delta\phi\sim$ a few tens of degrees cannot be ruled out, the observations do not support a sudden 90 degree rotation of the position angle at the jet break, which is a signature of top-hat axisymmetric jets. Given the sparsity of polarisation data points, a single-peaked polarisation curve remains consistent with the observations. The evolution of both the polarisation degree and position angle can be explained by structured jets, whether axisymmetric or not. If a gradual rotation of the position angle were confirmed, it would indicate a structured, non-axisymmetric jet.
@@ 5.2 GRB 091018
Figure 9: Sample fit for GRB091018 (red solid line). Top-hat jet model parameters: $e=0.15$, $\xi=0.25$, $\chi=-1.48$ and $P_{0}=0.15$. Optical polarimetry data from (Wiersema et al., 2012).
@@
Linear polarisation monitoring with VLT (Wiersema et al., 2012) revealed a polarisation curve inconsistent with a single-peaked profile, and instead favouring a two-peak structure. In top-hat jet models, both axisymmetric and non-axisymmetric, the second peak tends to have a higher amplitude than the first. Although some data points support this trend, significant scatter suggests the presence of a deep, short dip within the second peak. The PPA remains stable and nearly constant during the first peak, but shifts to a higher overall value in the second peak, with significant scatter around the 90-degree rotation. These variations in both the polarisation degree and position angle during the second peak are not fully explained even by non-axisymmetric jets (i.e., elliptical jet heads). As noted by Wiersema et al. 2012, if the scatter in the second peak is caused by an additional component such as a bright patch within the jet or microlensing effects that partially cancel the polarisation signals, the presence of a two-peak profile, combined with a nearly constant PPA in the first peak, suggests either a top-hat jet or a view of the core of a structured jet. In this scenario, the jet head (or core) must have a small eccentricity or be observed from a line of sight located near its semi-major or semi-minor axis. Additionally, the shallow post-break decay indices may imply that the jet core is encased by an outer region with a shallow angular profile.
@@ 5.3 GRB 121024A
Figure 10: Sample fit of GRB121024A (red solid line). Top-hat jet model parameters: $e=0.65$, $\xi=0.35$, $\chi=\pi/2$ and $P_{0}=0.22$. Optical polarimetry data from (Wiersema et al., 2014).
A long burst with $T_{90}=69$ sec was detected by Swift, with a redshift of $z=2.298$ determined shortly afterwards. The X-ray light curve follows a three-segment power-law evolution, with the final break at $t\sim 10$ hours coinciding with a break in the optical light curve (Wiersema et al., 2014). The monochromatic nature of this break suggests a jet break origin. However, the post-break decay indices remain relatively shallow, at 1.67 in the X-rays and 1.25 in the optical. VLT polarimetry measurements taken 3-6 hours after the burst show linear polarisation degrees ranging from 3-5%, with a stable position angle (see Figure 10). However, observations from the following night (after the jet break) reveal a significant shift in the position angle, consistent with a 90 degree rotation (Wiersema et al., 2014). These polarisation properties also support the similar jet structure interpretations as for GRB 091018, although the polarisation degree shows significant scatter (with larger uncertainties) after the break (the top panel), and our model curve does not reproduce these data well in this case either.
@@ 5.4 GRB 210610B
Figure 11: Sample fit of GRB210610B (red solid line). Top-hat jet model parameters: $e=0.7$, $\xi=0.35$, $\chi=-1.1$ and $P_{0}=0.2$. Optical polarimetry data from (Agüí Fernández et al., 2024).
A long-duration burst at $z=1.134$ was detected by Fermi ($T_{90}=69$ sec) and Swift ($T_{90}=55$ sec). The optical and X-ray light curves are modelled using a broken power-law. The optical light curve remains initially flat until the break at $\sim 7.8$ hours, after which it steepens to a decay index of 1.85. The X-ray light curve appears to undergo a simultaneous break and shares the same post break decay index, though its pre-break decay is faster than that of the optical counterpart (Agüí Fernández et al., 2024). Polarimetric observations were conducted with the Calar Alto Telescope and the VLT. The first measurements, taken at around 2.9 hours, revealed a relatively high polarisation degree of $\sim 4.5\%$ as shown in Figure 11. During the second observation period (5.8–6.7 hours), which is close to the break time, the polarisation level dropped significantly to 0.18–0.6%. However, after the break, the polarisation degree increased again to $\sim 2\%$ (30.4–30.6 hours). Interestingly, the PPA rotates by 54 degrees between the pre- and post-break measurements (the bottom panel), suggesting a significant shift that deviates notably from the 90 degree rotation expected for axisymmetric jets.
This event is a strong candidate for modelling with a non-axisymmetric jet. Here, we assume top-hat, non-axisymmetric jets for our analysis. While a structured, non-axisymmetric jet could also explain the event if the line of sight intersects the jet core, the constraint on the jet head’s eccentricity remain similar. Given the large deviation of $|\Delta\phi-\pi/2|=36$ degrees, and the distinct trough structure $R_{pd}\geq 10$, we performed 5000 random samplings. For a lower eccentricity ($e=0.3$), most cases (87%) yield $R_{pd}>10$. However, virtually none satisfy $|\Delta\phi-\pi/2|>$ 36 degrees. For $e=0.75$, the fraction of cases meeting the first, second and both conditions are 64%, 35% and 21%, respectively. The lines of sight that satisfy both conditions are characterised by intermediate azimuthal angles: $0.4<\chi<1.5$ and a relatively central region: $\xi<0.7$. These findings suggest this event could be explained within a non-axisymmetric jet framework. However, there are caveats in the jet-break modelling. 1) The flat optical light curve before the break suggests significant energy injection into the forward shock at early times. Since ejecta from the central engine can carry large-scale magnetic fields (Mundell et al., 2013) unlike forward shocks, which generates highly tangled magnetic fields, the polarisation signals before the break may be significantly influenced by emission from refreshed shocks (energy injection). 2) The post-break decay index of the optical afterglow ($1.85\pm 0.04$) is close to, but still shallower than, the expected value ($\sim 2$) for the standard fireball model (Agüí Fernández et al., 2024).
@@ 6 Conclusions
We have investigated the afterglow polarisation signals around a jet break for non-axisymmetric jets. For jets with top-hat luminosity density distributions, the polarisation degree curves exhibit two peaks in general, even when the jet head is elliptical. However, a complete polarisation drop between the peaks is generally absent. Unlike axisymmetric jets, where the polarisation position angle (PPA) undergoes a sudden jump at the trough, the PPA in non-axisymmetric jets evolves smoothly. Additionally, the change in PPA between the two peaks can deviate significantly from 90 degrees. In our random line of sight sampling, 50% of the cases show deviations of more than 18 degrees, and 10% exceed 31 degrees for $e=0.6$.
@@
As a simple toy model for a structured jet, we considered a case where the luminosity density in the jet’s outer region follows a power-law profile. While this simplified model should captures the key features of polarisation signals in structured jets, the angular dependence of luminosity density in a more realistic scenario arises from the lateral energy distribution - where the jet core has a higher energy per solid angle, and the outer regions gradually contain less energy per solid angle. This also induces the angular variations in the Lorentz factor, with the core having a higher Lorentz factor and progressively decreasing toward the outer regions. As a result, the visible region is no longer circularly symmetric, toward the jet centre, the relativistic beaming is stronger and luminosity contribution is larger (e.g., Figure 2 in Beniamini et al. 2022). The impact of this deformed visible region will be explored in a future paper.
When constructing the polarisation degree light curves shown in Figures 2, 6 -11, we assumed the forward shock dynamics $\Gamma\propto t^{-3/8}$ for the ISM. In the wind medium, the curve shapes are slightly modified due to different forward shock dynamics $\Gamma\propto t^{-1/4}$. However, the overall morphology (e.g. the presence of two peaks or a single peak) remains unchanged and the discussion on $\Delta\phi$ should be similar.
If the jet head has a highly eccentric shape, it can be characterised by two angular scales: $\theta_{j}$ (the opening angle along the semi-major axis) and $\theta_{j2}$ (along the semi-minor axis). As long as $\Gamma>1/\theta_{j2}$, the flux follows the typical blast wave emission evolution, decaying approximately as $t^{-1}$. When $\Gamma\propto t^{-3/8}$ drops below this first critical threshold, photons begin to be emitted over a larger solid angle, increasing from $\theta_{j}\theta_{j2}$ to $\Gamma^{-1}\theta_{j}$, causing the flux to decay as $t^{-11/8}$. As $\Gamma$ continues to decrease and falls below the 2nd threshold $1/\theta_{j}$, the emission solid angle expands further to $\Gamma^{-2}$, leading to a steeper decay of $t^{-7/4}$. Since these breaks arise from geometrical effects, they are expected to be monochromatic, similar to the usual jet break. The two jet break times $t_{j}$ and $t_{j2}$, with $t_{j2}<t_{j}$, satisfy the relation
@@
Although our case study section focused on optical afterglow observations, the jet break is a geometrical effect, meaning our results are applicable across different wavelengths (e.g., X-ray, radio). The number of events with measured radio polarisation has recently increased (Urata et al. 2019; Corsi et al. 2018; Laskar et al. 2019; Urata et al. 2023). Significant advancement in X-ray polarimetry in recent years (e.g., Astrosat, IXPE, POLAR, XPoSat) suggest that future X-ray satellites may be capable of detecting and analysing polarisation signals days after GRBs. Since both X-ray and optical bands are expected to lie above the typical frequency of the forward shock emission around jet breaks, the relativistic limb-brightening effect should be pronounced in both bands (Granot et al., 1999).
We adopted a conventional scenario in which magnetic fields are amplified by plasma kinetic instabilities at the shock, resulting in a coherence length scale comparable to the plasma skin depth scale (Medvedev & Loeb, 1999). Recently, large-scale turbulent magnetic fields driven by magnetohydrodynamic instabilities have also been proposed (e.g., Kuwata et al. 2023). In this scenario, the coherence length scale is comparable to the thickness of the blast wave. In the isotropic field case, both the polarisation degree and the PPA change randomly and continuously over time, resembling the behaviour described in the classical patch shell model (e.g., Gruzinov & Waxman 1999; Nakar & Oren 2004). By incorporating magnetic field anisotropy and the observer’s viewing angle, this model can also account for a variety of polarisation behaviour in afterglows. Notably, in this framework, the degree of radio polarisation can exceed that in the optical band (Kuwata et al., 2024). Simultaneous polarimetric observations across multiple wavelengths will be crucial for constraining the magnetic field amplification mechanism and improving our understanding of polarisation signals in afterglows.
@@ Acknowledgements
We thank the anonymous referee for their constructive comments and suggestions, which have helped to improve the manuscript. We are also grateful to Gavin P. Lamb for useful discussions. This research was supported by the Science and Technology Facilities Council (STFC) grants.